# Classical Simulations of Low Magic Quantum Dynamics

### v1

@@ IV.2 Space Complexity
We now turn to the space complexity in our new simulation algorithm. This is the number of entries, or bits and complex numbers, required to specify the state. The stabilizer state is represented by a tableau of size $(2L+1)\times(2L+1)$. Each stabilizer generator $\sigma_{l}$ is represented by a $2L$ binary vector, as defined in Section IV. Each phase $\lambda_{l}$ is a complex number. Each logical Pauli operator $\sigma_{l}$ is specified by an $L$-component vector with entries in $\{0,1,2,3\}$, labeling $I,X,Y$, and $Z$, respectively. These considerations mean that
@@
We present the number of entries $\overline{\mathcal{N}}$ as a function of $L$ in Fig. 4(a), under $p_{m,X}\in[0.6,0.9]$ at $p_{T}=1/L$, which sits in the area law magic phase of Fig. 6 We provide a comparison of $\overline{\mathcal{N}}=2^{L}$ statevector methods (dashed black line) with the LRSD (solid blue lines). We find that $\overline{\mathcal{N}}$ exhibits power law growth with some exponent $\alpha$, $\overline{\mathcal{N}}=\beta+L^{\alpha}$, with $\alpha\approx 2$. This is much slower than the exponential growth of statevector methods.
Figure 5: (a) Phase diagram of the purification transition in the single-pair all-to-all model with $X$-basis only measurements as a function of measurement probability $p_{m,X}$ and probability of $T$-gates, $p_{T}$. The part of the phase diagram labeled mixed is the mixed phase of the purification transition, where $\overline{S_{Q}}\sim e^{t/\tau(L)}$ and $\overline{\tau(L)}$ grows exponentially in $L$. The part of the phase diagram labeled pure is the pure phase of the purification transition, where $\overline{\tau(L)}$ is the area law in $L$. $\overline{\phantom{X}}$ denotes averaging over circuits. The critical measurement probability separating the mixed phase and pure phase is given by the critical purification (cp) rate $p_{m,X}^{cp}=0.26(1)$, which is weakly dependent on $p_{T}$. The points represent the numerically calculated locations of the purification transition, where $\tau(L)$ grows logarithmically in $L$. The black dashed line is the phase boundary deduced by interpolation. (b) Purification time of $\overline{S_{Q}}$ as a function of system size $L$ at $p_{T}=0$ for different measurement probabilities $p_{m,X}$ from $0.18$ to $0.36$ in the steps of 0.02. The critical measurement probability is $p_{m,X}^{cp}=0.26(1)$ with the dynamical exponent $z_{p}=0.22(2)$ (see Eq. (23)). At $p_{m,X}^{cp}$, $\tau\sim L^{z_{p}}$, as indicated by the dashed line. (c) $\tau L^{-z_{p}}$ as a function of measurement probability $p_{m,X}$ for $L\in\{16,32,64\}$. The inset shows the data collapse with the estimated critical exponent $\nu_{p}=0.45(5)$. The ensemble size is 25000 circuits for each value of $p_{m,X}$ and $L$ in all three panels.
@@ V.1 Purification Transition
The dynamical purification phase transition can be efficiently probed by studying how the system preserves entanglement over time with a single reference qubit. After preparing $|\psi_{o}\rangle=|+x\rangle^{\otimes L}$, the reference qubit is entangled to a randomly selected system qubit to form a Bell pair. The entangling operation is followed by a scrambling stage, which consists of $\sqrt{10}L$ random two-qubit Clifford gates applied to random system qubit pairs. We choose $\sqrt{10}L$ gates because the system’s all-to-all entanglement, Eq. 79 in Appendix B.2, is saturated at this point. This ensures that the ancilla is maximally entangled with all the qubits in the system, and not just a single qubit. This reduces the finite-size effects. We evolve the circuit up to a final time $t_{f}=2L^{2}$ in the single-pair all-to-all model with $X$-basis measurements. This is shown in 2(a). This timescale is chosen to ensure that the ancilla entanglement entropy has entered its long-time regime, where it decays exponentially with time for all system sizes and measurement rates. The decay of $\overline{S}_{Q}$ with $t$, for $0<t<2L^{2}$ is shown in 14 at $p_{m,X}=p_{m,X}^{cp}$.
When we use X-basis measurements, Eq. 20 no longer applies. Recall that we used Eq. 20 in Sec. IV.1 to argue that the $T$-gates constitute a single layer of single-qubit unitaries at the end of the Clifford circuit. The $T$ gates are commuted to the end of the circuit using the fact that they are diagonal in the $Z$ basis and commute with both the $CZ$ gates and $Z$-basis measurements. A final layer of single-qubit unitaries cannot alter the entanglement of a pure state $|\psi\rangle$. In this case, $|\psi\rangle$ is a stabilizer state, which has a degenerate entanglement spectrum. However, since Eq. 20 no longer applies, the argument at the beginning of Section IV.1 is invalid. The $T$-gates are therefore interspersed throughout the circuit, which means the entanglement is no longer the same as the stabilizer state. Instead, the final state is a generic non-stabilizer state without a degenerate entanglement spectrum. We must therefore use the LRSD rather than the stabilizer formalism to simulate the state and calculate the entanglement.
@@ V.2 Magic Transition
Figure 6: (a) Phase diagram of the magic and entanglement transitions in the single-pair all-to-all model with $Z$-basis only measurements as a function of the measurement probability $p_{m,Z}$ and $T$-gate rate $p_{T}=\eta/L^{\beta}$. The numerical results are shown for $\beta=1$ and $\eta\in\{0.25,\cdots,4.0\}.$ The green points show the numerically extracted locations of the magic transition. The entanglement transition is degenerate in the rate of $T$-gates, and was located using stabilizer simulations for $p_{T}=0.$ For low measurement rates and low $T$-gate rates, we find volume-law entanglement and area-law magic (phase I). For high measurement rates and low $T$-gate rates, we find area-law entanglement and area-law magic (phase II). For low measurement rates and high $T$-gate rates, we find volume-law entanglement and power-law magic (phase III). For high measurement rates and high $T$-gate rates, we find area-law entanglement and power-law magic (phase IV). (b) Scaling of the stabilizer nullity $\mathcal$$\overline${M}}, as defined in Eq. (15), with $L$. The $\overline{\phantom{X}}$ denotes averaging over circuits. The measurement probability is $p_{m,Z}=0.5$ and $1.6\leq\beta\leq 2.3$ and $\eta=1$. Blue curves denote the (sub)-extensive phase, the black line denotes logarithmic criticality, and the red lines denote the area law phase. Data is shown in the single-pair all-to-all circuit model with $Z$-basis measurements. (c) $\mathcal$$\overline${M}} divided by $\mathrm{log}(L)$ as a function of $\beta$ for different system sizes.
Here, we analyze a magic transition in the single-pair all-to-all model with only Z-basis measurements. Following the argument around Eq. 20, $Z$-basis measurements constrain the placement of $T$-gates in a structured way. This leads to a density matrix representation that is amenable to Bell sampling, as discussed in Appendix A.3. We then apply the algorithms developed in Appendix A.3 to efficiently compute the magic. The phase transition is between a (sub)-extensive and area law scaling of magic controlled by the rate of measurements $p_{m,Z}$ and probability of $T$-gates $p_{T}$. To find a phase transition in the magic, we study states that interpolate between states produced by Clifford and non-Clifford circuits in a controlled way. We set the probability to sample $T$ gates as
@@
Previous work in a bricklayer geometry has shown that if the density of non‑Clifford $T$ gates is kept constant as the system size $L$ grows, $\beta=0$, the total magic follows a volume law for any measurement rate $p<1$ [17]. Magic is additive and can be produced locally, whereas entanglement can only be generated by gates whose support straddles the boundary between two subsystems. Because each site receives a fixed injection of magic from a finite gate density, the total magic therefore scales extensively with $L$ for $\beta=0$. For $\beta=1$, $\eta$ represents average number of $T$ gates per time step, independent of $L$. The resulting density is $p_{T}=\eta/L^{\beta}$; choosing $\beta=1$ drives $p_{T}\to 0$ in the thermodynamic limit while still providing a finite reservoir of magic. We therefore focus on this $\beta=1$ protocol and discuss other scalings where appropriate.
@@
with the constraint that $\gamma>0$ corresponds to (sub)-extensive scaling, and $\gamma\leq 0$ to area-law behavior. If the best fit yields $\gamma\leq 0$, we classify the data as area-law; otherwise, we interpret it as (sub)-extensive. Details of the fitting procedure are provided in Appendix C. The values of magic for the trajectories plotted in Fig.6 empirically indicate a (sub)-extensive/log/area-law scaling of the magic with $L$. This is consistent with the scaling behavior of the magic, which was previously studied in bricklayer random Clifford circuits with a vanishing density of $T$-gates [17]. At the critical point $p=p_{c}$, we therefore invoke logarithmic scaling, i.e., $\mathcal{M}(L)=a+b\log L$ or
@@
In Fig. 6(a), we identify four distinct phases in the magic and entanglement. The phases are area-law magic, volume-law entanglement (Phase I), area-law magic, area-law entanglement (Phase II), power-law magic, volume-law entanglement (Phase II), and area-law entanglement, power-law magic (Phase IV). In this model, the entanglement and purification transitions coincide. The exact exponent $\beta$ of the scaling plays a crucial role. We show in Fig. 6(a) and (b) that, for sufficiently small $\beta$, the magic will scale trivially as (sub)-extensive for all values of $p_{m,Z}.$ Conversely, for sufficiently large $\beta$, the magic will scale as area law. For intermediate values of $\beta$ and for a fixed measurement rate, the magic will go from area to power-law scaling with increasing $\eta.$
@@ VI Discussion and conclusion
In this work, we introduced a simulation method based on the LRSD that stores and evolves the full density matrix of a near-stabilizer state. This capability allows access to quantities nonlinear in the state, such as entanglement. Using this approach, we identified—in the limit of $Z$-basis–only measurements—a MIPT in magic that is distinct from the MIPT in entanglement.
@@
For the $Z$-basis measurement model, we find four distinct phases in entanglement and magic. Phase II begins generating sub-extensive entanglement and magic—resources essential for universal quantum computation—while Phases II and IV remain low-entanglement and can be efficiently simulated with MPS. In the area-law magic regime (Phases III and IV), LRSD can further explore entanglement MIPTs in various circuit geometries and probe dynamical features such as operator growth and information scrambling via out-of-time-order correlators.
@@
We leave the possibility of how to evolve the LRSD under other non-Clifford resources to future work. Future work could also use this simulator to characterize the entanglement and magic MIPTs in other circuit geometries under $T$-gates. There may be a nontrivial flow of the universality class of the entanglement MIPT, for example, in other circuit geometries and/or injection of non-Clifford resources.
@@ Acknowledgements.
This work was partially supported by the Army Research Office Grant No. W911NF-23-1-0144 (K.A.  and J.H.P.), the US-ONR grant No. N00014-23-1-2357 (H.P. and J.H.P.). M.J.G. acknowledges support from the Defense Advanced Research Projects Agency (DARPA) under Agreement HR00112490357 and the NSF QLCI award OMA2120757. This work was performed in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452 (J.H.P.) and at the Kavli Institute for Theoretical Physics (KITP), which is supported by grant NSF PHY-2309135 (M.J.G., J.H.P.)
@@ A.1.1 Clifford Unitaries
Equation 29 gives the update rule for a Clifford unitary $U_{C}$. Because any Clifford maps Pauli operators onto Pauli operators by conjugation, each logical operator transforms as $\sigma_{l}\mapsto U_{C}\sigma_{l}U^{\dagger}$, and the stabilizer state transforms as $\rho_{S}\mapsto U_{C}\rho_{S}U_{C}^{\dagger}$. Consequently the LRSD retains its form after the update: the set of stabilizer generators $\{g_{i}\}$, the destabilizers $\{d_{i}\}$, and the logical Paulis $\{\sigma_{l}\}$ are merely mapped onto other Pauli operators by the conjugation, while the coefficients $\lambda_{l}$ remain unchanged. The LRSD is then updated as:

### v2

@@ IV.2 Space Complexity
We now turn to the space complexity in our new simulation algorithm. This is the number of entries, or bits and complex numbers, required to specify the state. The stabilizer state is represented by a tableau of size $(2L+1)\times(2L+1)$. Each Pauli operator $\sigma_{l}$ is represented by a $2L$ binary vector, as defined in Section IV. Each phase $\lambda_{l}$ is a complex number. Each logical Pauli operator $\sigma_{l}$ is specified by an $L$-component vector with entries in $\{0,1,2,3\}$, labeling $I,X,Y$, and $Z$, respectively. These considerations mean that
@@
We present the number of entries $\overline{\mathcal{N}}$ as a function of $L$ in Fig. 4(a), under $p_{m,X}\in[0.6,0.9]$ at $p_{T}=1/L$, which sits in the area law magic phase of Fig. 6 We provide a comparison of $\overline{\mathcal{N}}=2^{L}$ statevector methods (dashed black line) with the LRSD (solid blue lines). We find that $\overline{\mathcal{N}}$ exhibits power law growth with some exponent $\alpha$, $\overline{\mathcal{N}}=\beta+L^{\alpha}$, with $\alpha\approx 2$. This is much slower than the exponential growth of statevector methods. In the CAMPS representation, the computational cost remains polynomial for Clifford circuits and can often remain efficient when a Clifford circuit is “doped” with up to $\mathcal{O}(L)$ $T$-gates, particularly when these are distributed favorably across the system. In this regime, the disentangling procedure can frequently absorb the non-Clifford action without a significant increase in bond dimension [20]. Beyond this point—when the number of $T$-gates exceeds $\sim L$ or their placement generates unavoidable entanglement—the bond dimension typically grows rapidly, leading to exponential cost. For sufficiently sparse $T$-gate densities, such as rates scaling as $p_{t}\sim 1/L^{2}$, the simulations remain tractable, and similar behavior is expected for small prefactors $\eta$ in the scaling form $p_{t}\sim\eta/L$. Within the area-law magic phase, the LRSD algorithm exhibits robust power-law scaling with system size $L$ (Fig. 4), which plausibly provides an upper bound on the scaling with $T$-gate count at fixed system size.
Figure 5: (a) Phase diagram of the purification transition in the single-pair all-to-all model with $X$-basis only measurements as a function of measurement probability $p_{m,X}$ and probability of $T$-gates, $p_{T}$. The part of the phase diagram labeled mixed is the mixed phase of the purification transition, where $\overline{S_{Q}}\sim e^{t/\tau(L)}$ and $\overline{\tau(L)}$ grows exponentially in $L$. The part of the phase diagram labeled pure is the pure phase of the purification transition, where $\overline{\tau(L)}$ is the area law in $L$. $\overline{\cdots}$ denotes averaging over circuit trajectories. The critical measurement probability separating the mixed phase and pure phase is given by the critical purification (cp) rate $p_{m,X}^{cp}=0.26(1)$, which is weakly dependent on $p_{T}$. The points represent the numerically calculated locations of the purification transition, where $\tau(L)$ grows logarithmically in $L$. The black dashed line is the phase boundary deduced by interpolation. (b) Purification time of $\overline{S_{Q}}$ as a function of system size $L$ at $p_{T}=0$ for different measurement probabilities $p_{m,X}$ from $0.18$ to $0.36$ in the steps of 0.02. The critical measurement probability is $p_{m,X}^{cp}=0.26(1)$ with the dynamical exponent $z_{p}=0.22(2)$ (see Eq. (21)). At $p_{m,X}^{cp}$, $\tau\sim L^{z_{p}}$, as indicated by the dashed line. (c) $\tau L^{-z_{p}}$ as a function of measurement probability $p_{m,X}$ for $L\in\{32,64,128\}$. The inset shows the data collapse with the estimated critical exponent $\nu_{p}=0.45(5)$. The ensemble size is 25000 circuits for each value of $p_{m,X}$ and $L$ in all three panels.
@@ V.1 Purification Transition
The dynamical purification phase transition can be efficiently probed by studying how the system preserves entanglement over time with a single reference qubit. After preparing $|\psi_{o}\rangle=|+x\rangle^{\otimes L}$, the reference qubit is entangled to a randomly selected system qubit to form a Bell pair. The entangling operation is followed by a scrambling stage, which consists of $\sqrt{10}L$ random two-qubit Clifford gates applied to random system qubit pairs. We choose $\sqrt{10}L$ gates because the system’s all-to-all entanglement, Eq. (77) in Appendix B.2, is saturated at this point. This ensures that the ancilla is maximally entangled with all the qubits in the system, and not just a single qubit. This reduces the finite-size effects. We evolve the circuit up to a final time $t_{f}=2L^{2}$ in the single-pair all-to-all model with $X$-basis measurements. This is shown in Fig. 2(a). This timescale is chosen to ensure that the ancilla entanglement entropy has entered its long-time regime, where it decays exponentially with time for all system sizes and measurement rates. The decay of $\overline{S}_{Q}$ with $t$, for $0<t<2L^{2}$ is shown in Fig. 14 at $p_{m,X}=p_{m,X}^{cp}$.
When we use X-basis measurements, Eq. (18) no longer applies. Recall that we used Eq. (18) in Sec.IV.1 to argue that the $T$-gates constitute a single layer of single-qubit unitaries at the end of the Clifford circuit. The $T$ gates are commuted to the end of the circuit using the fact that they are diagonal in the $Z$ basis and commute with both the $CZ$ gates and $Z$-basis measurements. A final layer of single-qubit unitaries cannot alter the entanglement of a pure state $|\psi\rangle$. In this case, $|\psi\rangle$ is a stabilizer state, which has a degenerate entanglement spectrum. However, since Eq.  (18) no longer applies, the argument at the beginning of Section IV.1 is invalid. The $T$-gates are therefore interspersed throughout the circuit, which means the entanglement is no longer the same as the stabilizer state. Instead, the final state is a generic non-stabilizer state without a degenerate entanglement spectrum. We must therefore use the LRSD rather than the stabilizer formalism to simulate the state and calculate the entanglement.
@@ V.2 Magic Transition
Figure 6: (a) Phase diagram of the magic and entanglement transitions in the single-pair all-to-all model with $Z$-basis only measurements as a function of the measurement probability $p_{m,Z}$ and $T$-gate rate $p_{T}=\eta/L^{\beta}$. The numerical results are shown for $\eta=12$ and $p_{m,Z}\in\{0.0,\cdots,1.0\}.$ The green points show the numerically extracted locations of the magic transition. The entanglement transition is degenerate in the rate of $T$-gates, and was located using stabilizer simulations for $p_{T}=0.$ For low measurement rates and low $T$-gate rates, we find volume-law entanglement and area-law magic (phase I). For high measurement rates and low $T$-gate rates, we find area-law entanglement and area-law magic (phase II). For low measurement rates and high $T$-gate rates, we find volume-law entanglement and power-law magic (phase III). For high measurement rates and high $T$-gate rates, we find area-law entanglement and power-law magic (phase IV). (b) Scaling of the stabilizer nullity $\mathcal$$\overline${M}}, as defined in Eq. (13), with $L$. The $\overline{\cdots}$ denotes averaging over circuit trajectories. The measurement probability is $p_{m,Z}=0.5$ and $0.6\leq\beta\leq 1.3$ and $\eta=12$. Blue curves denote the (sub)-extensive phase, the black line denotes logarithmic criticality, and the red lines denote the area law phase. Data is shown in the single-pair all-to-all circuit model with $Z$-basis measurements. (c) $\mathcal$$\overline${M}} divided by $\mathrm{log}(L)$ as a function of $\beta$ for different system sizes.
Here, we analyze a magic transition in the single-pair all-to-all model with only Z-basis measurements. Following the argument around Eq. (18), $Z$-basis measurements constrain the placement of $T$-gates in a structured way. This leads to a density matrix representation that is amenable to Bell sampling, as discussed in Appendix A.3. We then apply the algorithms developed in Appendix A.3 to efficiently compute the magic. The phase transition is between a (sub)-extensive and area law scaling of magic controlled by the rate of measurements $p_{m,Z}$ and probability of $T$-gates $p_{T}$. To find a phase transition in the magic, we study states that interpolate between states produced by Clifford and non-Clifford circuits in a controlled way. We set the probability to sample $T$ gates as
@@
Previous work in a bricklayer geometry has shown that if the density of non‑Clifford $T$ gates is kept constant as the system size $L$ grows, $\beta=0$, the total magic follows a volume law for any measurement rate $p<1$ [19]. Magic is additive and can be produced locally, whereas entanglement can only be generated by gates whose support straddles the boundary between two subsystems. Because each site receives a fixed injection of magic from a finite gate density, the total magic therefore scales extensively with $L$ for $\beta=0$. For $\beta=1$, $\eta$ represents average number of $T$ gates per time step, independent of $L$. The resulting density is $p_{T}=\eta/L^{\beta}$; choosing $\beta=1$ drives $p_{T}\to 0$ in the thermodynamic limit while still providing a finite reservoir of magic.
@@
with the constraint that $\gamma>0$ corresponds to (sub)-extensive scaling, and $\gamma\leq 0$ to area-law behavior. If the best fit yields $\gamma\leq 0$, we classify the data as area-law; otherwise, we interpret it as (sub)-extensive. Details of the fitting procedure are provided in Appendix C. The values of magic for the trajectories plotted in Fig. 6 empirically indicate a (sub)-extensive/log/area-law scaling of the magic with $L$. This is consistent with the scaling behavior of the magic, which was previously studied in bricklayer random Clifford circuits with a vanishing density of $T$-gates [19]. At the critical point $\beta=\beta_{c}$, we therefore invoke logarithmic scaling, i.e., $\mathcal{M}(L)=a+b\log L$ or
@@
In Fig. 6(a), we identify four distinct phases in the magic and entanglement. The phases are area-law magic, volume-law entanglement (Phase I), area-law magic, area-law entanglement (Phase II), power-law magic, volume-law entanglement (Phase III), and area-law entanglement, power-law magic (Phase IV). In this model, the entanglement and purification transitions coincide. The exact exponent $\beta$ of the scaling plays a crucial role. Consider parameters $p_{m}$ and $\eta$ in regions III or IV. For some finite $\beta$, these points lie in the volume-law magic phase, while for $\beta=\infty$, no magic is present. We show numerically in Fig. 6(a), that for sufficiently low $p_{m}$, $\beta_{c}>1$ and decreases to $\beta_{c}=1$ for increasing $p_{m}$. This suggests that the $\beta=\infty$ “zero-magic” fixed point is stable. Conversely, for $p_{m}$ and $\eta$ in regions I or II, the system is in the area-law magic phase for some finite $\beta$, and is in the volume-law magic phase at $\beta=0$. For sufficiently high $p_{m}$ and low $\eta$, it is likely that a critical point $\beta_{c}<1$ exists.
We leave to future work as to whether, as $p_{m}\to 0$ and $\eta\to\infty$, the critical value $\beta_{c}$ also diverges. In other words, whether the $\beta=\infty$ “zero-magic” fixed point is unstable, implying that the system resides in the magic phase for any positive $\beta$. Similarly, it remains an open question whether, for sufficiently large $p_{m}$ and small $\eta$, $\beta$ tends to zero, placing the system in the area-law magic phase for any finite $\eta$.
@@ VI Discussion and conclusion
In this work, we introduced a simulation method based on the LRSD that stores and evolves the full density matrix of a near-stabilizer state. This capability allows access to quantities nonlinear in the state, such as entanglement. Using this approach, we identified, in the limit of $Z$-basis-only measurements, a MIPT in magic that is distinct from the MIPT in entanglement.
@@
For the $Z$-basis measurement model, we find four distinct phases in entanglement and magic. Phase III begins generating sub-extensive entanglement and magic—resources essential for universal quantum computation—while Phases II and IV remain low-entanglement and can be efficiently simulated with MPS. In the area-law magic regime (Phases I and II), the LRSD can further explore entanglement MIPTs in various circuit geometries and probe dynamical features such as operator growth and information scrambling via out-of-time-order correlators.
@@
Data Availability
The data used in this work is available from the authors upon request.
@@ Acknowledgements.
We thank David Huse, Grace Sommers, and Bryan Clark for useful discussions. This work was partially supported by the Army Research Office Grant No. W911NF-23-1-0144 (K.A.  and J.H.P.), the US-ONR grant No. N00014-23-1-2357 (H.P. and J.H.P.). K.A. and M.J.G. acknowledge support from the Defense Advanced Research Projects Agency (DARPA) under Agreement HR00112490357 and the NSF QLCI award OMA2120757. We are grateful to the OSG Consortium [39, 40, 41, 42] for providing us computational resources to run simulations. This work was performed in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452 (J.H.P.) and at the Kavli Institute for Theoretical Physics (KITP), which is supported by grant NSF PHY-2309135 (M.J.G., J.H.P.)
@@ A.1.1 Clifford Unitaries
Equation (27) gives the update rule for a Clifford unitary $U_{C}$. Because any Clifford maps Pauli operators onto Pauli operators by conjugation, each logical operator transforms as $\sigma_{l}\mapsto U_{C}\sigma_{l}U^{\dagger}$, and the stabilizer state transforms as $\rho_{S}\mapsto U_{C}\rho_{S}U_{C}^{\dagger}$. Consequently the LRSD retains its form after the update: the set of stabilizer generators $\{g_{i}\}$, the destabilizers $\{d_{i}\}$, and the logical Paulis $\{\sigma_{l}\}$ are merely mapped onto other Pauli operators by the conjugation, while the coefficients $\lambda_{l}$ remain unchanged. The LRSD is then updated as: